Aquariums require a delicate balance of dissolved gases. A clown fish requires at least 7.0 mg/L O2 to survive, but greater than 30.0 ppm of CO2 is toxic. You want to setup a 20.0 L aquarium tank at 25.0 °C, where only the atmosphere is responsible for dissolved gases (you’ve just added the fish, so ignore O2 consumption and CO2 production).

PO2 = 0.21 atm PCO2 = 0.0039 atm

20. What is the minimum O2 concentration where the fish can survive?

A. 2.2 x 10-4 M   B. 0.219 M  C. 4.57 M  D. 0.011 M
Answer #1

Sol.

Minimum O2 Concentration required for fish to survive in mg/L = 7 mg/L   

Minimum O2 Concentration required for fish to survive in g/L = 7 mg/L × 1g / 1000 mg = 0.007 g/L  

Also , Molar Mass of O2 = 32 g/mol

So , Minimum O2 Concentration required for fish to survive in mol/L ( M )  

= 0.007 g/L / 32 g/mol  

= 2.2 × 10-4 mol/L ( M )  

Therefore , answer is   option A , 2.2 × 10-4M